If there are 356 beans in a 16 ounce bag, how many beans would there be in a 20 ounce bag?
Answer by Alan3354(69443)

If there are 356 beans in a 16 ounce bag, how many beans would there be in a 20 ounce bag?
---------------
356/16 = 22.25 beans/ounce
22.25*20 = 445 beans
===============
356:16::n:20
16n = 356*20
n = 445 beans
============
20 = 16*1.25
356*1.25 = 445 beans
===========
I'm going with 445 beans